### v1

# Advancing Reference-free Evaluation of Video Captions with Factual Analysis

Abstract
Video captions offer concise snapshots of actors, objects, and actions within a video, serving as valuable assets for applications such as question answering and event localization. However, acquiring human annotations for video captions is costly or even impractical, especially when dealing with diverse video domains. Existing models trained on supervised datasets face challenges in evaluating performance across different domains due to the reliance on reference-based evaluation protocols, which necessitate ground truth captions. This assumption is unrealistic for evaluating videos in the wild. To address these limitations, we propose a reference-free evaluation framework that does not require ground truth captions, focusing on factual grounding to ensure accurate assessment of caption quality. We introduce VC-Inspector, a novel caption quality evaluator that is both reference-free and factually grounded. Utilizing large language models, we generate pseudo captions of varying quality based on supervised data, which are subsequently used to train a multimodal model (i.e., Qwen2.5-VL) as the evaluator. Our approach demonstrates superior alignment with human judgments on the VATEX-Eval dataset, outperforming existing methods. The performance also generalizes to image caption datasets, Flickr8K-Expert and Flickr8K-CF, when viewing images as 1-frame videos. Overall, VC-Inspector offers a scalable and generalizable solution for evaluating the factual accuracy of video captions, paving the way for more effective and objective assessment methodologies in diverse video domains.
1 Introduction
Figure 1: Existing reference-free metrics like EMScore [31] often fail to detect factual inaccuracies and lack a consistent scoring scale. VC-Inspector addresses these limitations by providing factually grounded, interpretable evaluations with explanations.
Video captioning plays a pivotal role in bridging visual content with natural language, offering concise descriptions of salient entities, actions, and interactions within a video. These captions are instrumental in enabling a wide range of downstream applications, including video-based question answering [27], event localization [28, 12], and content retrieval [8]. To ensure the quality and reliability of captions, robust evaluation metrics are vital for advancing the field of video-language research. However, current captioning models, predominantly trained on supervised datasets [14, 41], rely heavily on reference-based evaluation protocols [26, 3, 19, 33, 39] that compare generated captions against human-written ground truth. While effective in controlled settings, these metrics face significant limitations in real-world scenarios. First, generating such references is prohibitively labor-intensive and costly, often becoming unavailable or infeasible to obtain when scaling across diverse video domains with varying visual and contextual characteristics. Second, these metrics demonstrate a limited ability to understand the semantic meanings of the captions and often fail to recognize valid alternatives.
The challenges of assessing open-domain or in-the-wild videos necessitate the development of evaluation protocols that do not rely on reference captions. However, this remains underexplored within the video research community. Current reference-free metrics [31, 30] typically measure the visual-language semantic alignment using pretrained multimodal embeddings [29], with the comprehension of caption content restricted by the context length of the underlying text encoder. In addition, these metrics frequently overlook factual inaccuracies in the candidate captions, as shown in Figure 1. Furthermore, they are primarily image-based solutions, making them suboptimal for video content.
To address these limitations, we propose a novel reference-free evaluation framework, VC-Inspector, that eliminates the dependency on human-annotated captions. Our approach emphasizes factual grounding, ensuring that the evaluation of caption quality is based on the alignment between the caption and the actual video content, rather than comparison to a predefined reference. We build VC-Inspector atop a lightweight large multimodal model (LMM), equipping the model with the ability to judge the quality of video captions based on factual accuracy, through instruction tuning. However, while supervised video caption datasets are already scarce, captions of imperfect quality are even harder to obtain. We tackle this challenge by introducing a novel data generation pipeline powered by a large language model (LLM), creating pseudo captions of varying quality by altering the factual elements (i.e., objects and actions) in the ground truth captions from ActivityNet [14]. This assembles ActivityNet-FG-It, comprising 44K data points for instruction tuning. Figure 2 illustrates the data generation and training pipeline of VC-Inspector. Unlike previous metrics that offer only a single quality score, our approach also requires the model to explain the rationale behind its scoring, making the evaluator more interpretable.
We start with evaluating VC-Inspector on the ActivityNet-FG-Eval and YouCook2-FG-Eval datasets, where the labels are synthetically generated with the same data generation pipeline as the instruction tuning dataset. Results demonstrate that the quality estimates are consistent across datasets, and VC-Inspector is versatile for different visual content and caption lengths. We further evaluate the correlation between the quality scores given by VC-Inspector and human judgments using the VATEX-Eval dataset [31], a standard benchmark for video captioning metric evaluation. The proposed evaluator stands out compared to the previous reference-free metrics, even surpassing most of the reference-based metrics. Finally, we extend the evaluation to image caption datasets, Flickr8K-Expert and Flickr8K-CF [10], showing the generalization of the proposed metric.
Overall, this work makes the following contributions:
  • We present one training dataset and two evaluation datasets consisting of ground truth captions, synthetic candidate captions, and pseudo quality scores, generated using a novel data generation pipeline.
  • We propose VC-Inspector, an open-source reference-free evaluator featuring factual grounding, with a text context length of 32K tokens and support for long videos.
  • Our evaluator outperforms existing reference-free metrics for video captions, achieving higher correlations with human evaluations and generalizing across visual domains.
2 Related Works
Traditional text-based evaluation can be extended to video captioning by discarding the video and treating the reference caption as ground truth. Alternatively, image-caption metrics can be adapted by sampling images from the video and aggregating image-caption scores. We review these metrics and their limitations below:
Text-only metrics based on references.
Several rule-based metrics have been developed to evaluate generated texts, including widely used methods such as BLEU [26], METEOR [3], ROUGE [19], and CIDEr [33]. While these metrics perform reasonably well in controlled settings, they primarily focus on syntactic structure and fail to capture the semantic meaning of the generated text. To address this limitation, the SPICE series [1, 18] attempts to embed semantic understanding by parsing both reference and candidate captions into objects, attributes, and their relationships. However, these metrics still struggle to account for semantically similar or identical words expressed differently. This persistent challenge has motivated the development of embedding-based evaluation methods, such as BERTScore [40] and its extended version [37], which consider the intrinsic variance between multiple ground truth captions. More recently, CLAIR [5] explores an LLM-as-a-Judge approach for image caption evaluation, demonstrating stronger correlation with human judgments than the above metrics. However, these metrics all rely exclusively on reference captions for evaluation and do not incorporate the visual input that the captions are meant to describe.
Image-augmented metrics.
To address the limitation of text-only metrics, various approaches have been explored that cross-reference between visual input and captions. VIFIDEL [24] computes semantic similarity by matching object names extracted from the image with those mentioned in the candidate caption. However, objects are discrete representations limited to fixed categories and cannot capture the motion dynamics that are particularly informative for videos. Visual-language models (VLMs), by contrast, provide continuous representations that offer richer semantic alignment between visual and textual modalities. ViLBERTScore [16] extends BERTScore by leveraging the pretrained ViLBERT [23] models, but still requires comparison to reference captions. UMIC [15] and CLIP-based metrics [9, 13, 7, 30, 34, 38, 17, 31, 21] relax this constraint by employing contrastive learning to directly model the semantic alignment between the images and captions. Although promising for short captions, these embedding-based methods are constrained by the context length limitations of their text encoders. Recently, the strong contextual reasoning and instruction following capability of LLMs have been leveraged for evaluation across diverse tasks [22]. Building on this trend, Maeda et al. [25] employ an LLM to compare the candidate caption to the visual context extracted with a VLM. However, these methods only consider static images and do not model the temporal dynamics of video, resulting in suboptimal performance on video caption evaluation.
Figure 2: (left) We present a data generation pipeline designed to systematically create synthetic video captions with diverse quality scores, along with explanations for the assigned scores. (right) This dataset was subsequently used for instruction tuning the VC-Inspector.
In summary, text-based methods require costly, high-quality reference captions, while image-based extensions are suboptimal for video content. These challenges highlight the growing demand for reference-free video caption evaluation tailored for video inputs. However, it remains an underexplored topic in the community, compared to text-based and image-caption evaluation. EMScore [31] was an early attempt that supports evaluating video captions without a reference. Although it considers both frame-level and video-level embeddings, these embeddings are still obtained from an image-based encoder [29], and the text encoder is limited by a short context length. PAC-S [30] and FactVC [21] augment EMScore with positive and negative data synthesis, respectively. While they share some similarity to this work, they consider only a single level of corruption, with binary (positive/negative) differentiation, whereas our method incorporates captions with varying degrees of quality, enabling a more nuanced evaluation. In addition, these metrics produce a single scalar score without offering explanations of their judgments, posing challenges for interpreting the quality evaluation. More recently, G-VEval [32] extended the G-Eval [22] approach to video by stitching together three frames from each video. However, it is unclear whether this image-based prompting generalizes to longer or more dynamic videos. Moreover, the dependence on proprietary models such as GPT-4o limits its scalability and practicality for widespread use. In contrast, VC-Inspector finetunes a lightweight, open-source LMM to produce both evaluation scores and reasoning, offering a more interpretable and scalable solution.
Ground Truth Candidate Caption Rouge-L SPICE Soft-SPICE
The man is feeding a cat on the sofa in the living room The man is feeding a lion on the sofa in the living room 92.31 75.00 78.57
The girl is dancing in the room The girl is sleeping in the room 85.71 66.67 88.19
Table 1: Text-based metrics rely on ground truth for evaluation and often fail to detect factual errors in objects/actions (bolded).
3 Video Caption Quality Estimation
In this section, we provide a concise overview of existing metrics used for evaluating video captions, identify their shortcomings, and outline our objectives.
3.1 Overview
[…]
3.2 Toward Factually Grounded Reference-free Evaluation
In this paper, we aim to develop a factually grounded and reference-free video caption quality evaluator that can generate scores on an absolute scale, and is versatile for various domains and text lengths. We argue that an evaluation protocol is only reliable if it is factually grounded, meaning that the quality assessment must accurately reflect the correctness of object entities and actions in the caption with respect to the video. For example, the evaluator should assign a lower score to a candidate caption where an object is missing or an action is wrong in the video than a correct one, and an even worse score for a caption that has more errors. However, it remains unclear whether existing metrics can effectively capture these subtle changes in factual elements. To investigate this, we probe existing metrics with candidate captions that contain incorrect objects or/and actions. Table 1 and Figure 1 illustrate such cases for reference-based and reference-free metrics, respectively. These results reveal that current metrics are insensitive to these factual errors. Even when captions are semantically distinct (e.g., “cat" vs. “lion" and “dancing" vs. “sleeping"), high scores are assigned due to substantial syntactic overlap.
Motivated by these observations, we propose VC-Inspector, a novel factually grounded and reference-free model for inspecting the video caption quality. We leverage the recent advents of large multimodal models (LMMs) to handle long-context text reasoning and generalized video feature extraction, with the hypothesis that their established efficacy on visual-language joint reasoning can be applied to this task. While evaluation metrics are preferred to be lightweight, we employ the 3B/7B version of Qwen2.5-VL as a model-based evaluator and equip the model with the factual grounding ability through instruction tuning, where the model receives a video $V$ and a candidate caption $\hat{X}$, and generates an integer score ranging from 1 to 5. We also request the model to provide an explanation for the assigned score in the output, similar to [32], which makes the evaluator more interpretable and can serve as a supervision signal for factual grounding. The next section will cover how we systematically generate captions with incorrect factual elements, the corresponding quality estimates, and the explanations.
4 VC-Inspector: Instruction Tuning for Factually Grounded Evaluation
In video captioning, two primary factors that can degrade caption quality are: (1) the inclusion of objects not present in the video, and (2) the introduction of actions not depicted in the video. Our preliminary study indicates that existing metrics fall short in capturing these factual elements in the caption. To address this, we propose to empower LMMs with this ability through instruction tuning. However, it is nontrivial to acquire annotated samples for training. Supervised video caption datasets are already scarce due to the high cost of labeling, and these datasets generally only include correct captions, lacking those of lower quality. To obtain a considerable number of samples with diverse quality, we propose a systematic way to create a synthetic video caption dataset leveraging the internal knowledge of large language models (LLMs).
Figure 3: Data generation pipeline to create a synthetic dataset for training VC-Inspector. While both “talking” and “holding” were identified as actions, only “holding” was sampled for replacement in the synthetic dataset.
4.1 Data Generation
To create captions with incorrect elements, we employed Llama-3.3-70B-Instruct to alter the objects and actions in a ground truth caption from a supervised video caption dataset. Specifically, we created the ActivityNet-FG-It dataset for instruction tuning utilizing the ActivityNet [14] training set. The data generation pipeline is summarized in Figure 3.
[…]
After generating a pseudo caption, it is critical to assign it a reasonable quality score that is intuitive for humans so that it can better mirror humans’ expectations. To achieve this, one could utilize a visual-language model (VLM) to estimate this score; however, VLMs are not robust enough to produce a factually grounded score in a specified range without calibration, and the scoring quality is confined by their capabilities. Instead, we employ a deterministic scoring mechanism based on factual grounding, described as follows,
[…]
$=1-\frac{|{\cal R}|}{|{\cal O}|+|{\cal A}|}.$ (1)
Since $\cal{R}\subseteq{\cal O}\cup{\cal A}$ and $|{\cal O}\cup{\cal A}|=|{\cal O}|+|{\cal A}|$, the score will reside within the range of 0 and 1. It also guarantees that a caption with more incorrect objects/actions receives a lower score. We then convert the score to an integer in the 1-5 range for better interpretability during model training.
We repeated the aforementioned process to create 10 pseudo captions per ground truth caption, resulting in a total of 374K pseudo captions derived from 37,396 video-caption pairs. The prompts used to guide the generation process are detailed in the supplementary. The randomized replacement of objects and actions ensures coverage across the full range of the possible scores, as the number of replacements directly influences the semantic deviation from the original caption. However, it also naturally yields a skewed and non-uniform score distribution. To mitigate the potential bias during training, we applied a balanced sampling strategy, resulting in a refined subset of approximately 218K pseudo captions with uniform representation across five score categories. However, due to computational constraints, training with the full 218K instances would require multiple weeks of runtime. Therefore, we further sampled a 44K (8.8K for each label) caption subset from the balanced dataset for instruction tuning, which we refer to as ActivityNet-FG-It. This subset preserves category balance and maintains the diversity and structure of the original data, while offering a tractable size for experimentation ($\sim 32$ GPU hours using A100).
4.2 Training
[…]
where we format the explanation $E$ in text using the information collected in the data creation process, i.e., the list of changed objects and actions, detailed in the supplementary. This can serve as extra supervision for the model to implicitly learn factual grounding and provide interpretable reasoning for this model-based evaluator at test time. During training, the model is optimized with the language modeling loss, like in other instruction tuning work [20].
5 Experiments
5.1 Experimental Settings
Datasets.
We train VC-Inspector with the proposed instruction tuning dataset, ActivityNet-FG-It, which comprises 44K video-caption pairs, along with their quality scores and explanations. To evaluate whether VC-Inspector can generalize to other videos and visual domains, we further follow the same data generation pipeline in section 4.1 to create two evaluation datasets, ActivityNet-FG-Eval and YouCook2-FG-Eval, according to the ActivityNet [14] test set and YouCook2 [41] validation set. However, these are not truly labeled by humans. To evaluate whether the quality scores given by VC-Inspector are aligned with human evaluators, we utilized the widely adopted VATEX-EVAL dataset, specifically designed for evaluating video caption evaluation quality. It contains 6 captions with varying levels of quality per video, each rated by three human evaluators on a scale of 1 to 5. Since some videos become unavailable on YouTube, we collect the remaining subset of 2,590 videos and the corresponding 15,540 candidate captions. Unless otherwise specified, all experiments were evaluated on the same dataset to ensure a fair comparison. We further extend our evaluation to image caption datasets, Flickr8K-Expert and Flickr8K-CF [10], by viewing images as single-frame videos to test the generalization of the proposed metric. The former contains 17K image-caption pairs rated by three human experts on a scale of 1 to 4, while the latter collects binary quality assessments for 48K image-caption pairs from crowd sources.
Baselines.
The baselines are organized into three categories: i) Language-based metrics: The evaluation in this category solely relies on the reference texts without considering the visual input. Representative metrics are BLEU [26], ROUGE [19], METEOR [3], CIDEr [33], BERTScore [39], SPICE [1], SPICE-Factual [18], Soft-SPICE  [18] and CLAIR [6]. ii) Image-augmented metrics: These approaches incorporate images as references (alongside reference captions), for better capturing the semantic alignment between the visual input and the candidate caption, e.g., CLIPScore [9], EMScore [31], PAC-S [30], FactVC [21], and G-VEVAL [32]. They usually support both reference-free and reference-based settings, where the former compares the candidate caption directly to the video without reference captions, while the latter uses both visual and textual references. We report results for both settings for completeness, but our primary focus is on the reference-free setting, which is more practical for real-world, in-the-wild videos. iii) Video-based metrics: These methods employ a video encoder to incorporate full video sequences as references. To the best of our knowledge, no existing metric in the literature falls into this category. Therefore, we adapt CLIPScore [9] to the recent advent of ViCLIP [35] as a stronger baseline, ViCLIPScore. Additionally, we compare VC-Inspector against the base model it builds upon, the vanilla Qwen2.5-VL model, to highlight the benefit of fine-tuning for evaluative reasoning.
Metrics.
In the caption evaluation literature, the effectiveness of evaluation metrics is typically assessed by measuring their correlation to human judgments. Following prior work [31, 32], we report both Kendall’s correlation ($\tau_{b}$) and Spearman’s rank correlation ($\rho$).
Implementation details.
VC-Inspector is developed for two model sizes, 3B and 7B, initialized from their corresponding Qwen2.5-VL pretrained weights. In all experiments, we train the model on 4 NVIDIA-A100 GPUs with a global batch size of 128 and a learning rate of 1e-4. We set both alpha and rank to 32 for the low-rank adaptation with a dropout rate of 0.05. During inference, we use a temperature of 0.0 for reproducibility.
5.2 Generalization of Quality Estimation
In this section, we evaluate the consistency of quality estimates using synthetic data, and validate the effectiveness of VC-Inspector by measuring its alignment with human judgments on the video caption evaluation dataset, VATEX-Eval. Given the limited availability of video caption datasets, we treat images as single-frame videos and extend our evaluation to image caption datasets, Flickr8K-Expert and Flickr8K-CF, to assess the generalization.
Evaluation with the synthetic data.
ActivityNet-FG-Eval YouCook2-FG-Eval
Metric $\tau_{b}$ $\rho$ $\tau_{b}$ $\rho$
EMScore [31] 28.94 40.77 20.21 29.24
CLIPScore [9] 28.10 39.65 18.00 26.14
Qwen2.5-VL-3B [2] 37.91 47.80 37.16 47.17
VC-Inspector-3B 49.53 62.01 44.29 55.31
Table 2: Correlation scores on the synthetic ActivityNet-FG-Eval [4] and YouCook2-FG-Eval [14] datasets. The best score is bolded.
Table 2 presents the correlation between metric scores and ground truth annotations in the ActivityNet-FG-Eval and YouCook2-FG-Eval datasets, both of which contain pseudo captions with varying degrees of factual inaccuracies. VC-Inspector, specifically finetuned for factual grounding, consistently outperforms baselines in differentiating incorrect captions. Notably, although it is only trained on ActivityNet-FG-It, its quality estimation generalizes effectively across visual domains. This suggests that the proposed data generation pipeline produces stable and reliable quality scores, rather than noisy outputs that fail to converge.
Evaluation on VATEX-Eval.
Table 5.2 presents our primary results: human correlation scores on the VATEX-EVAL [31] dataset. Baselines are grouped by reference type as outlined in section 5.1, and the three column-sections correspond to the No Reference, 1-Reference, and 9-Reference settings, respectively. Language-based metrics, which rely heavily on textual references, are not applicable to our target No Reference setting. Therefore, we focus our comparisons on multimodal methods that incorporate visual inputs. VC-Inspector consistently outperforms all evaluated metrics in the reference-free setting, particularly those based on CLIP embeddings, including CLIPScore, EMScore, FactVC, and PAC-S. When adapting CLIPScore to the recently introduced ViCLIP [35] (which adopts a video encoder), we observe a considerable gain over those image-CLIP-based approaches. Despite ViCLIPScore being a stronger baseline, it still underperforms relatively to VC-Inspector and its underlying model, as the context length of the ViCLIP model (i.e., 32) is much shorter than the 32K context length of ours, which supports a more flexible and comprehensive caption evaluation. We further compare VC-Inspector with G-VEval [32], a recent LMM-based evaluation method. However, G-VEval is based on GPT-4o, a proprietary, paid model with an estimated 200B parameters. In contrast, VC-Inspector is open-source, lightweight, and reproducible, with configurations available at 3B and 7B parameters depending on system requirements. While having a compact size, our 7B model achieves the highest correlation to human evaluations and could potentially enable on-the-fly quality estimation during training, making it a viable reward model in Reinforcement Learning (RL) applications. Although our primary focus is to compare with the reference-free metrics, it is noteworthy that VC-Inspector even surpasses the performance of most reference-based metrics.
Evaluation on Flickr8K.
No Reference 1-Reference 9-References
Metric $\tau_{b}$ $\rho$ $\tau_{b}$ $\rho$ $\tau_{b}$ $\rho$
[…]
BLEU_1 [26] - - 12.65 16.52 28.70 36.88
BLEU_4 [26] - - 12.44 36.28 22.76 29.56
ROUGE-L [19] - - 12.94 16.89 23.94 31.06
METEOR [3] - - 16.68 21.80 27.64 35.76
CIDEr [33] - - 17.62 23.02 27.92 36.18
BERTScore [39] - - 15.24 19.82 25.05 32.37
SPICE [1] - - 14.80 18.78 27.41 35.40
SPICE-factual [18] - - 13.59 17.04 26.05 35.58
Soft-SPICE [18] - - 21.25 27.61 36.31 46.41
CLAIR [6] - - 36.00 - 34.80 -
[…]
CLIPScore [9] 22.33 29.09 27.39 35.49 35.21 45.28
EMScore [31] 22.88 29.79 28.63 37.05 36.66 47.00
FactVC [21] 22.79 29.69 28.78 37.22 36.18 46.33
PAC-S [30] 25.10 - 32.60 - 31.40 -
G-VEval [32] 39.40 - 44.90 - 48.10 -
[…]
ViCLIPScore [9, 35] 30.92 39.86 - - - -
Qwen2.5-VL-3B [2] 31.29 36.43 - - - -
Qwen2.5-VL-7B [2] 34.70 39.40 - - - -
VC-Inspector-3B 37.99 42.45 - - - -
VC-Inspector-7B 42.58 45.99 - - - -
Table 3: Human correlation scores on the VATEX-EVAL [31] dataset. indicates results reported from [32] due to GPT-4 licensing issue. The best score for each column section is bolded. Please note that this work focuses on the No Reference setting. Our best model has outperformed all other models in this setting, while remaining competitive with metrics that rely on references.
Metric Flickr8K-Expert Flickr8K-CF
[…]
BLEU_1 [26] 32.20 17.90
BLEU_4 [26] 30.60 16.90
ROUGE [19] 32.10 19.90
METEOR [3] 41.50 22.20
CIDEr [33] 43.60 24.60
SPICE [1] 51.70 24.40
BERTScore [39] - 22.80
CLIPScore [9] 52.60 36.40
PAC-S [30] 55.50 37.60
[…]
CLIPScore [9] 51.10 34.40
PAC-S [30] 53.90 36.00
VC-Inspector-3B 59.86 39.00
VC-Inspector-7B 63.43 45.97
Table 4: Correlation score ($\tau_{b}$) with human judgments on Flickr8k-Expert and Flickr8k-CF [10] dataset. The overall best scores are bolded and best of each section is underlined. Our model has outperformed even the reference-based methods.
To expand our evaluation beyond the limited availability of video caption datasets, we treat images as an extreme case of short videos with a single frame. Table 5.2 reports the human correlation score on two widely adopted benchmarks, Flickr8K-Expert and Flickr8K-CF. The former requires a more fine-grained differentiation among the captions, with ratings distributed across 4 levels, whereas the latter employs binary judgments. VC-Inspector is instruction-tuned with captions that exhibit varying degrees of factual inaccuracies, enabling nuanced evaluation of caption quality. This allows the model to perform effectively across both benchmarks, despite their differing rating scheme. In these evaluations, VC-Inspector remains the best-performing method under the reference-free setting, substantially narrowing the gap to inter-human correlation ($\tau_{b}\approx 73$[1], and even outperforming several reference-based metrics. These results demonstrate the strong generalization capability of VC-Inspector across visual domains and video lengths.
5.3 Ablation and Analysis
We conduct ablation studies on VC-Inspector components on the VATEX-Eval benchmark:
Data synthesis strategies.
Since both objects and actions are informative factual elements for a video, grounding the evaluator in factual understanding requires instructing the model to identify errors in these components. To this end, during the data generation process (outlined in section 4.1), we systematically altered both elements in ground truth captions to create pseudo captions for training. In this section, we ablate the impact of modifying these elements in the ActivityNet training set by evaluating three variants: i) Changing objects only, ii) Changing actions only, and iii) Changing both. As shown in Table 5, all variants achieve strong alignment with human ratings compared to EMScore. However, the variant that alters both objects and actions yields the best performance. These results highlight the importance of both factual elements, object and action, in capturing the context of video content for caption evaluation. They also demonstrate the robustness and generalization of the proposed paradigm across different factual errors.
Role of explanations.
Figure 4: Visual examples from ActivityNet-FG-Eval (top) and VATEX-Eval (others). VC-Inspector produces quality assessments highly consistent with ground truth scores, along with explanatory insights into factual errors (highlighted in red).
Metric Data Synthesis $\tau_{b}$ $\rho$
EMScore [31] n/a 22.88 29.79
VC-Inspector-3B Change objects only 36.40 41.20
Change actions only 33.23 39.63
Change both (Ours) 37.99 42.45
Table 5: Ablation study on synthetic data generation strategies. Human correlation scores ($\tau_{b}$, $\rho$) are reported on the VATEX-Eval.
During our data generation process, we systematically synthesize pseudo captions along with their corresponding quality estimates. This process yields an informative “side product": explanations that indicate where factual errors are located within the candidate captions. We leverage these explanations as an auxiliary supervision signal during training, enabling the model to implicitly learn factual grounding, which has been shown to be effective compared to the variant without explanations in Table 6. Not only are explanations useful in training, but they also enhance the interpretability of the model-based metric, providing transparency of why the scores are assigned to given captions. Further evaluation of the explanation quality is provided in the supplementary.
Metric stability.
To evaluate the stability of our metric, following prior work [36], we computed the Pearson correlation between scores obtained across two runs. We find a perfect correlation of 1.0 (as expected with temperature 0.0), confirming that VC-Inspector produces stable quality estimation across multiple runs.
VC-Inspector-3B $\tau_{b}$ $\rho$
Without Explanations 34.29 38.18
With Explanations 37.99 42.45
Table 6: Impact of explanations on the model performance. In the “Without Explanation" setting, we have trained the model only on the score, removing the pseudo explanations.
Computational efficiency.
We assess the computational efficiency by comparing the average runtime per video clip. EMScore [31], ViCLIPScore [35], and VC-Inspector require 0.42, 0.34, and 0.30 seconds per clip, respectively, on a single A100 GPU. VC-Inspector demonstrates superior efficiency compared to existing methods.
5.4 Qualitative Results
Figure 4 presents visual examples of VC-Inspector outputs on the ActivityNet-FG-Eval (top) and VATEX-Eval (middle, bottom) datasets. Without relying on reference captions, the model evaluates candidate captions based on the associated video content and produces quality scores that closely mirror human judgments. The explanation pinpoints the factual inaccuracies, such as incorrect objects and/or actions in the candidate captions, and assigns factually grounded scores accordingly. Additional qualitative results are provided in the supplementary.
6 Conclusion and Discussion
This work addresses the challenge of evaluating video captions across diverse domains without relying on human-annotated reference captions. We identified the limitations of existing metrics, and proposed a novel reference-free and factually grounded evaluation framework, VC-Inspector, based on LMMs. By delving into the factual analysis of captions, we empower the LMM with the factual grounding ability through instruction tuning, where we systematically create pseudo captions of diverse quality using an LLM. Experimental results across multiple domains demonstrate that VC-Inspector achieves high consistency with human evaluators, and outperforms existing metrics on detecting factual errors. Its versatility and interpretability make it a practical tool for evaluating the factual accuracy of video captions in real-world settings.
While our current focus is on factual errors, particularly object and action substitutions, the proposed paradigm can be potentially extended to simulate other types of errors (e.g., attribute replacements), as supported by our ablations on synthesizing only one type of error. Additionally, we recognize the importance of evaluating temporal coherence and narrative structure, which remain underexplored. Future work should expand this paradigm to simulate and detect a broader range of errors, enabling more comprehensive and context-aware evaluation strategies for video captions.
Appendix
The appendix is organized as follows. Section A lists the detailed prompts used in the proposed data generation pipeline, model fine-tuning, and explanation evaluation. Section B provides the evaluation of explanations generated by VC-Inspector, and section C presents additional visual examples with candidate captions of diverse quality.
Appendix A Prompts
The data generation prompts are reported on the following blocks:
  • Extract object – Prompt A
  • Extract action – Prompt A
  • Find similar object – Prompt A
  • Find similar action – Prompt A
  • Substitute object or action – Prompt A
The fine-tuning prompt and the prompt for evaluating the generated explanations are reported in Prompt A and Prompt A, respectively.
Prompt A.1: Prompt to extract object from a caption ### Instruction: Given the input text, generate a list of objects in the caption in the format of [ "Object1", "Object2", …]. Don’t include any verbs. ONLY REPLY THE ANSWER.
### Input: {{caption}}
### Output
Prompt A.2: Prompt to extract actions from a caption ### Instruction: Given the input text, generate a list of actions in the caption in the format of ["Action1", "Action2", …]. ONLY REPLY THE ANSWER.
### Input: {{caption}}
### Output
Prompt A.3: Prompt to find similar object given a object ### Instruction: Find the parent class of the given object and generate one of its child class that has a different meaning but shares the same parent. The new class cannot be a synonym or similar terms to the original object. It can be an antonym or any co-hyponym. For example, generate "dog" for "cat". ONLY REPLY THE NEW CLASS.
### Input: {{object}}
### Output:
Prompt A.4: Prompt to find similar action given an action ### Instruction: Find a different action that the subject can perform that has a different meaning than the input action. The new action cannot be a synonym or similar terms to the original action. For example, generate "put into" for "take out of". ONLY REPLY THE NEW ACTION.
### Input: {{action}}
### Output:
Prompt A.5: Prompt to substitute object or action given the caption and new object or actions ### Instruction: Substitute {{old_obj_act}} in {{caption}} as {{new_obj_act}} . Keep the answer in the same format as {{caption}} . ONLY REPLY THE ANSWER.
### Input: cap
### Output:
Prompt A.6: Fine-tuning prompt ### USER:
[[VIDEO]]
<caption> {{caption}} </caption> You are given a video and a caption describing the video content. Please rate the helpfulness, relevance, accuracy, level of details of the caption. The overall score should be on a scale of 1 to 5, where a higher score indicates better overall performance. Please first output a single line containing only one integer indicating the score. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias. STRICTLY FOLLOW THE FORMAT.
### ASSISTANT:
{{quality_score}}
The caption does not accurately capture the video content. For example, the actions ( {{wrong_act}} ) are incorrect / the objects ( {{wrong_obj}} ) are incorrect/the objects ( {{wrong_obj}} ) and actions ( {{wrong_act}} ) are incorrect.
Prompt A.7: Prompt for explanation evaluation [Context]
{{ground_truth_caption}}
[Caption]
{{caption_to_evaluate}}
[Groundtruth]
{{ground_truth_explanation}}
[End of Groundtruth]
[Assistant]
{{predicted_explanation}}
[End of Assistant]
[System]
We would like to request your feedback on the performance of an AI assistant in the response to the quality evaluation of the caption provided above with respect to a video. For your reference, the visual content in the video is represented with a few sentences describing the same video. You are also given a ground truth evaluation to that caption. Please rate the helpfulness, relevance, accuracy, level of details of the response by comparing to the ground truth and referring to the context information. Provide an overall score on a scale of 1 to 10, where a higher score indicates better overall performance.
Please first output a single line containing the score. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias.
Appendix B Evaluation on Explanation Quality
We further assess the quality of the explanations generated by VC-Inspector on two synthetic datasets, as the VATEX-Eval dataset does not offer ground truth explanations. Table 7 presents two metrics for this experiment. The BERT score measures the semantic similarity between the generated explanations and the pseudo ground truth explanations, according to pretrained textual embeddings. On the other hand, the LLM Score is derived by using the prompt detailed in section A to engage Llama-3-8B-Instruct in a LLM-as-a-judge approach, similar to [20]. We report the relative score assigned by the LLM that rescales the score received by the predicted explanations with respect to the score given to the ground truth explanations. Results show that the explanations provided by VC-Inspector, along with its quality assessment, align closely with the (pseudo) ground truth explanations obtained during the data generation process.
Dataset BERT Score LLM Score
ActivityNet-FG-Eval 0.79 93.11
YouCook2-FG-Eval 0.70 90.97
Table 7: Evaluation of the explanations generated by VC-Inspector-3B on two synthetic evaluation datasets.
Figure 5: Additional visual examples on VATEX-Eval. VC-Inspector identifies the incorrect objects (highlighted in red) and assigns scores aligned with human evaluators.
Figure 6: Additional visualization of VC-Inspector results on ActivityNet-FG-Eval videos, with candidate captions of diverse quality. Incorrect objects and actions are identified by VC-Inspector and labeled in red.
Appendix C Additional Visual Examples
Figure 5 presents additional visual examples of VC-Inspector applied to VATEX-Eval. Consistent with the results shown in Figure 4 of the main paper, VC-Inspector effectively identifies incorrect objects and actions in candidate captions, assigning quality scores that closely align with human judgments. In Figure 6, we showcase two video examples from ActivityNet-FG-Eval, each paired with candidate captions containing a progressively increasing number of factual inaccuracies. The model successfully detects these incorrect elements, provides detailed explanations, and yields scores that reflect the severity of the factual errors.